# Context-Sensitive Abstractions for Reinforcement Learning with Parameterized Actions

### v1

@@ B Test Environments
We evaluate in domains with long-horizons and sparse rewards i.e., agents receive a positive reward only upon reaching the goal state: (i) OfficeWorld (icarte2022reward; corazza2024expediting) (Fig. 4(a)): In this domain, the agent must navigate a cluttered indoor office environment to pick up mail and coffee and deliver them to designated office locations. The state space includes the agent’s $(x,y)$ position and two binary variables indicating whether it is carrying coffee or mail. The action space consists of four parameterized movement actions—one for each cardinal direction-with displacement values in the range $[0,0.5)$. The agent automatically picks up items when at their location and drops them off when at the target office location.
(ii) Pinball (roice2024new; rodriguez2024learning) (Fig. 4(b)): The agent controls a small ball in a physics-based arena and must guide it into a red hole, avoiding collisions with irregularly shaped obstacles. The ball is subject to dynamic physical forces, such as bouncing off obstacles and surface resistance. The action space includes five parameterized actions: four to increase or decrease velocity in the $x$ or $y$ direction, and one no-op action.
(iii) Multi-city transport (ma2021hierarchical; oswald2024large) (Fig. 4(d)): This domain models a complex, multi-city delivery problem. The agent navigates roads within cities and uses air transport to travel between them. The objective is to retrieve a package in one city and deliver it to a destination city. The environment includes three cities, each with an airport. The agent has five parameterized actions: up, down, left, right (each parameterized by distance), and a fly action (parameterized by the destination city), which can only be executed at airports.
(iv) Robot Soccer Goal  (bester2019multi) (Fig. 4(c)): The task involves an agent learning to kick a ball past a keeper. Three actions are available to the agent: kick-to(x,y), shoot-goal-left(y), and shoot-goal-right(y). It terminates if the ball enters the goal, is captured by the keeper, or leaves the play area.

### v2

@@ B Test Environments
We evaluate in domains with long-horizons and sparse rewards i.e., agents receive a positive reward only upon reaching the goal state: (i) OfficeWorld (Icarte et al. 2022; Corazza et al. 2024) (Fig. 4(a)): In this domain, the agent must navigate a cluttered indoor office environment to pick up mail and coffee and deliver them to designated office locations. The state space includes the agent’s $(x,y)$ position and two binary variables indicating whether it is carrying coffee or mail. The action space consists of four parameterized movement actions—one for each cardinal direction-with displacement values in the range $[0,0.5)$. The agent automatically picks up items when at their location and drops them off when at the target office location.
(ii) Pinball (Roice et al. 2024; Rodriguez-Sanchez and Konidaris 2024) (Fig. 4(b)): The agent controls a small ball in a physics-based arena and must guide it into a red hole, avoiding collisions with irregularly shaped obstacles. The ball is subject to dynamic physical forces, such as bouncing off obstacles and surface resistance. The action space includes five parameterized actions: four to increase or decrease velocity in the $x$ or $y$ direction, and one no-op action.
(iii) Multi-city transport (Ma et al. 2021; Oswald et al. 2024) (Fig. 4(d)): This domain models a complex, multi-city delivery problem. The agent navigates roads within cities and uses air transport to travel between them. The objective is to retrieve a package in one city and deliver it to a destination city. The environment includes three cities, each with an airport. The agent has five parameterized actions: up, down, left, right (each parameterized by distance), and a fly action (parameterized by the destination city), which can only be executed at airports.
(iv) Robot Soccer Goal  (Bester et al. 2019) (Fig. 4(c)): The task involves an agent learning to kick a ball past a keeper. Three actions are available to the agent: kick-to(x,y), shoot-goal-left(y), and shoot-goal-right(y). It terminates if the ball enters the goal, is captured by the keeper, or leaves the play area.